# Polarized and unpolarized gluon PDFs: generative machine learning applications for lattice QCD matrix elements at short distance and large momentum

### v1

@@ I Introduction
To achieve the goal of calculating momentum fraction, $x$-dependent nonperturbative structure functions and PDFs from the first-principles lattice QCD (LQCD) calculations, there have been several proposals, such as the path-integral formulation of the deep-inelastic scattering hadronic tensor [1, 2], the operator product expansion [3], current-current correlations [4], the Compton amplitude approach [5, 6], quasi-PDFs and large momentum effective theory (LaMET) [7, 8], lattice cross-sections [9, 10, 11], and pseudo-PDFs [12]. While LQCD calculations have made significant computational achievements in calculating PDFs (see the recent reviews [13, 14] and the references therein), one defining challenge is that the bilocal light-cone correlators that are necessary to determine the PDFs cannot be evaluated directly on the Euclidean lattice. Quasi-PDFs framework [7] circumvents this drawback by calculating equal-time Euclidean nonlocal matrix elements with hadron states at non-zero momentum, $p_{z}$. The corresponding quasi-PDFs can be matched to the light-cone PDFs when the hadron momentum is large, by applying the LaMET formalism [8]. On the other hand, the pseudo-PDFs framework uses short-distance factorization and can be perturbatively matched to the light-cone PDF. To implement the QCD short-distance factorization with controllable power corrections and because the renormalons in the renormalization of the LQCD matrix elements at large distances become significantly important [15], it is desirable to use only the short-distance matrix elements. However, one needs LQCD matrix elements at large Ioffe-time $\omega\equiv-z\cdot p$ [16, 17, 18] while keeping $z$ small and making the hadron boost $p_{z}$ along the $z$-direction as large as possible to facilitate the inverse problem in determining PDFs from a limited $\omega$-range data. The problem of using large-$z$ matrix elements has been found in a recent calculation [19] with the implementation of the 2-loop matching [20] as the higher-twist contribution can become significant for $z\gtrsim 0.35$ fm (see for other findings in [21, 22, 23]). An alternative proposal is to use the pseudo-PDFs ratio scheme [12] only at small $z$ in hybrid- and self-renormalization schemes [24, 25], treating short and long-distance scales separately as well as using model-dependent extrapolations of lattice correlation functions at large $\omega$ to facilitate the determination of PDFs within the LaMET framework.
@@
In this work, we train the generative machine learning (ML) algorithms using short-distance LQCD correlation functions and focus on the generation of the LQCD matrix elements at large $\omega$ which can be used to determine the polarized gluon $x\Delta g(x)$ and the unpolarized gluon $xg(x)$ PDFs. The principle behind training ML algorithms with short-distance correlation functions is based on the ability of the ML to learn effectively from the data points where higher-twist contamination is minimal. Therefore, it is expected that the generation of the renormalized matrix elements at large $\omega$ will also involve small contamination from higher-twist effects. Moreover, the ML algorithms predicting the LQCD matrix element at large $\omega$ for the larger values of $p_{z}$ will not only be extremely useful for handling the inverse problem of reconstructing the PDFs but also be particularly helpful in future calculation of quark and gluon PDFs within the LaMET fromalism [8].
@@
We note that there have been a few calculations of the unpolarized gluon PDFs in the nucleon, pion, and kaon [26, 27, 28, 29, 30, 31, 32, 33, 34, 35], as well as matrix element calculations toward the gluon helicity Ioffe-time distribution [36], and the first LQCD determination of the gluon helicity PDF [37]. In our work, we will show that determining the light-cone quark and gluon helicity PDFs require LQCD data in the limit of $p_{z}\to\infty$. We will demonstrate how physics-informed ML can help attain this limit and enable us to determine helicity PDFs, particularly the gluon helicity PDF. We will use published LQCD matrix elements of the unpolarized and polarized gluon correlation functions [38, 39] from Refs. [31, 36, 37]. The essential goals of this work are as follows:
@@
The paper is organized as follows. In Sec. II, we first discuss the necessity of extrapolating LQCD data outside the available range of $\omega$ and discuss how this extrapolation has been performed previously. This section reveals the risk of using model-dependent and overly constrained PDF ansatz and motivates the need for ML to generate data at large $\omega$ where LQCD data are not available. In Sec. III, we discuss the theoretical essence of LQCD data at large momentum and the prospect of generative ML models to overcome the above-mentioned problems. We discuss a prescription to eliminate the $p_{z}$-dependence in the lattice data of quark and gluon helicity correlation functions. In Secs. IV and V, we explore three different generative ML algorithms to determine the polarized and unpolarized gluon Ioffe-time distributions [18] beyond the reach of present-day LQCD calculations. In Secs. VI and VII, we determine $x\Delta g(x)$ and $xg(x)$ distributions using neural network on the ML-generated Ioffe time distributions and highlight some of the important implications of this work. Sec. VIII contains our concluding remarks and outlook.
@@
In general, LQCD calculation of matrix elements at hadron momentum $p_{z}>2~{}{\rm to}~{}3$ GeV becomes exponentially challenging due to poor signal-to-noise ratio. Specifically, achieving precise results for the gluonic matrix elements at a lighter pion mass and $p_{z}>2$ GeV will be very challenging in the near future LQCD calculations as it involves Wick contractions associated with the lattice calculations of disconnected insertions. The state-of-art LQCD calculations of the unpolarized gluonic matrix elements or the Ioffe-time distribution (ITD) have provided data up to $\omega_{\rm max}\approx 7$ [31, 29] or even smaller [32], beyond which the data is consistent with zero due to large noise. One way to facilitate the extraction of PDFs from the lattice data in a wide range of $\omega$, is to perform LQCD calculation at large spacelike separations at a fixed $p_{z}$. However, this poses the risk of uncontrollable power corrections, and the factorization of the LQCD matrix elements into nonpertubative PDF and perturbative coefficients might not be sensible. Additionally, at a fixed $p_{z}$, lattice data becomes much noisier as $z$ increases. Therefore, when one parameterizes the PDFs using some model-constrained functional forms and includes possible higher-twist contributions such as $\mathcal{O}(\Lambda_{\rm QCD}^{2}z^{2})$ or some higher order $\mathcal{O}(z^{n})$ terms in parameterizing the data, the noisy data may not allow for proper isolation of the higher-twist contributions from the leading twist-2 contribution (for example, see Fig. 5 in [40] at $z=0.75$ fm and $p_{z}\geq 1.67$ GeV).
@@
In addition, constrained by a limited range of $\omega$, the available lattice ITDs are inherently sensitive to only the first few moments, as shown in [41] and reiterated in [42, 43, 23]. Lattice calculations using model-dependent parametrization of PDFs utilize the correlation between the first two or three accessible moments to model the missing information beyond the available $\omega_{\rm max}$ and attempt to extract PDF in the full $x$ range. In addition, for currently available LQCD calculations in a limited $\omega$ range, these functional forms can be biased, leading to an unreliable $\chi^{2}/{\rm d.o.f.}$, and underestimation of uncertainties. For example, the parametrization $x^{\alpha}(1-x)^{\beta}$ for the LQCD determination of $xg(x)$ distribution led to a diverging PDF in [27] and a converging PDF in [31] at small $x$. However, none of these lattice ITDs reach the Regge region or have much sensitivity to the small-$x$ physics [43]. In [31], $\alpha\geq 0$ constraint was imposed in a Bayesian fit, motivated by a phenomenological analysis in [43]. Otherwise, it would have resulted in a diverging PDF as in [27]. A similar observation can be made for $xg(x)$ extractions in recent lattice calculations [26, 28, 29, 30, 32, 33] where all these lattice ITDs are limited in ITD the range of $\omega\lesssim 6$. It is therefore desirable to determine LQCD matrix elements at large $\omega$, while maintaining controllable power corrections. Additionally, it is important not to rely on simple ansatz fits of PDFs using data in a limited $\omega$ range as the resulting PDFs can depend heavily on the chosen functional forms, priors, and other strong constraints used in the fits. None of the existing LQCD data is sensitive to either very large $x$ or very small $x$ PDFs. It is not evident why a product of $x^{\alpha}$ (governed by the Regge behavior [44]) and $(1-x)^{\beta}$ (governed by the perturbative QCD counting rule [45, 46, 47]) should describe PDFs in the mid-$x$ region, where the lattice data is more sensitive to.
@@
As an illustration of the problem, by reconstructing the unpolarized gluon ITD from NNPDF distribution [48], it can be shown that the NNPDF ITD has significantly smaller uncertainty across the entire $\omega$ range compared to the lattice ITD presented in [31]. However, the large-$x$ PDF determined from the LQCD data has much smaller uncertainty than the NNPDF distribution. Although the Fourier transform of $\tilde{\omega}=\frac{1}{x}$ [10, 49] and precise lattice data at small $\omega$ can be helpful to constrain PDF in the large-$x$ region, such a dramatic reduction of the uncertainty in the large-$x$ PDFs is not expected. This possibility of underestimating the uncertainty in the lattice calculation of PDFs has been elaborated in Fig. 1, where the lattice and the phenomenological ITDs $\mathcal{I}_{g}(\omega,\mu)$ are chosen to be at $\mu=2$ GeV matching scale in the $\overline{\rm{MS}}$ scheme. Keeping in mind that this is not a one-to-one comparison because of the different types of observables used in the determination of PDFs, the reconstructed PDF from the lattice data using a Bayesian fit with the functional form of PDF $x^{\alpha}(1-x)^{\beta}$ (with imposed constraint $\alpha\geq 0$) in Fig. 1 has uncertainty smaller than the NNPDF determination in the whole $x\geq 0.7$ region. Therefore, it can be misleading to obtain a much more precise PDF from noisier lattice data in a limited $\omega$ range. The reason for this unrealistic precision in [31] stems from two constraints:
@@
  • Bayesian fits using Jacobi polynomial basis [50, 51, 52] with a constraint $\alpha\geq 0$.
A more direct comparison using the same lattice data with or without using constrained functional forms in the determination of gluon helicity PDFs has been illustrated in the lower panel in Fig. 1 (from Ref. [37]) which shows the severity of underestimating uncertainties when constrained PDF ansatz is used.
@@
Figure 1: Upper left panel: Ioffe-time distribution after the implementation of the perturbative matching kernel on the lattice reduced pseudo-ITD calculated for a $2$-parameter model fit, in the $\overline{\rm{MS}}$ renormalization scheme at 2 GeV in Ref. [31]. The $z=a-6a$ denotes the length of the field separation as a function of the lattice spacing $a$. The Ioffe-time distribution of the NNPDF distribution is normalized to $1$ at $\omega=0$ for a comparison with the LQCD matrix elements. Upper right panel: unpolarized gluon PDF (cyan band) extracted from the lattice data in Ref. [31] using the $2$-parameter model fit $x^{\alpha}(1-x)^{\beta}$. We compare our results to the gluon PDF extracted from the NNPDF3.1 global fit [53]. Lower panel: a similar example is given for the gluon helicity PDF from [37]. The cyan band is determined directly from the lattice QCD data without assuming any functional form for the PDF ansatz, while the other bands with smaller uncertainties result from different model fits used to extract the $x\Delta g(x)$ distribution. The 2-parameter fit is done using the fit-expression $x^{\alpha}(1-x)^{\beta}$, and subsequently by introducing additional parameters for 3 and 4 parameter fits, namely $\rho$ and $\gamma$, resulting in the fit-expression $x^{\alpha}(1-x)^{\beta}(1+\rho\sqrt{x}+\gamma x)$.
@@
Building on the findings presented in [37], this section briefly outlines the prescription for removing contamination terms in the Euclidean matrix elements that may obstruct the extraction of light-cone quark and gluon helicity PDFs.
To determine $\Delta g(x)$, we calculate matrix elements of the gluon field strength tensor $G_{\mu\nu}$ and its dual $\widetilde{G}_{\lambda\beta}=(1/2)\epsilon_{\lambda\beta\rho\gamma}G^{\rho\gamma}$ separated by a spatial Wilson line $W[z,0]$ [7, 39],
@@
where $z$ is the separation between the gluon fields, $p$ is the nucleon four-momentum, and $s$ is the nucleon polarization. The combination that can be used to determine the gluon helicity correlation with the least number of contamination terms is [39]
@@
$i,j=x,y$ being perpendicular to the nucleon boost in the $z$-direction, $p=\{p_{0},0_{\perp},p_{z}\}$. Utilizing the multiplicative renormalizability of the quasi-PDF matrix elements [54, 55, 56, 57], one can use the pseudo-PDFs ratio method [12] and obtain the renormalization group invariant reduced pseudo-ITD, $\Delta\mathfrak{{M}}$ [39]:
@@
where, $m_{p}$ is the nucleon mass, ${\cal M}_{00}(z,p_{z})\equiv[M_{0i;i0}(z,p_{z})+M_{ji;ij}(z,p_{z})]$ is the spin-averaged matrix element related to the unpolarized gluon correlation [38, 31] and the factor $1/Z_{\rm L}(z/a_{L})$ is determined in [39] to cancel the UV logarithmic vertex anomalous dimension of $\Delta{\cal M}_{00}$.
$\Delta\mathfrak{{M}}$ can be expressed in terms of invariant amplitudes, $\Delta\mathcal{M}_{sp}^{(+)}$ and $\Delta{\mathcal{M}}_{pp}$ [39],
@@
where the subscript $lc$ denotes taking the light-cone limit $z=z_{-}$ and applying the UV renormalization resulting in the dependence on the factorization scale $\mu$. Note that light-cone ITD in Eq. (5) does not contain the $m_{p}^{2}/p_{z}^{2}$ suppressed term as appeared in Eq. (4). An alternative expression of $\Delta\mathcal{M}_{00}(z,p_{z})$ shows that this matrix element is nonvanishing at $p_{z}=0$ and one can remove the $\mathcal{O}(\omega)$ contamination following the method describing in [36]. However, the residual higher-order contamination can become significant at large $\omega$. In order to eliminate the $(m_{p}^{2}/p_{z}^{2})\omega\Delta{\mathcal{M}}_{pp}$ contribution, we multiply Eq. (4) by the corresponding lattice squared-momentum $p_{k}^{2}$ and obtain [37]
@@
Following the derivation in [58], it can be shown that in a fixed gauge the nonlocal spatial matrix element in Eq. (2) reduces to $\Delta G\rightarrow\left(\vec{E}\times\vec{A_{\perp}}\right)^{3}$ in the infinite momentum frame, consistent with the findings in [59, 60], where $\vec{A}_{\perp}$ is the transverse part of the gauge potential.
As another example, a contamination term is also present in the LQCD matrix elements associated with the quark helicity PDF [40]. Following the same notations as in [40], the Euclidean matrix elements needed for the calculation of quark helicity PDF can be written as
@@
Now, the essence of the ML algorithms to utilize the relation in Eq. (7) can be understood as the following. The immediate challenge of implementing Eq. (7) is that the amplitudes $\Delta{\mathcal{M}}_{sp}^{(+)}$ and $\Delta{\mathcal{M}}_{pp}$ are nonperturbative functions and their functional forms are unknown. It has been shown in [37] that in the absence of a theoretically known functional form of these nonperturbative amplitudes, the moments expansion fails to describe the data, and one needs to resort to ML to circumvent the ignorance of the functional forms of these amplitudes. Therefore, to parameterize these unknown nonperturbative amplitudes, one can use some ML approaches constrained by Eq. (7) which we refer to as the “physics-informed” ML. Note that, as long as the equality in Eq. (7) holds to a good approximation, the lattice correlation functions will be dominated by the leading-twist contribution and can be used to extract the light-cone $x\Delta g(x)$. As a remarkable outcome, we will see in Sec. IV.1 that this relation is beneficial in exposing possible higher-twist contaminations in the gluon correlation functions at $z\gtrsim 0.36$ fm. Therefore, if ML approaches can be developed and trained on the lattice matrix elements at small $z$, where higher-twist contamination is minimal, and then used to generate the ITD at large $\omega$, the resulting extrapolated data may exhibit reduced higher-twist contamination even at larger values of $\omega$ for a fixed $p_{z}$. The ideal scenario is when ample and precise lattice data points are available, allowing ML to determine the correlations between different $z$ and $p_{z}$ datasets and estimate higher-twist contaminations. This will be a subject of future study with more precise LQCD matrix elements. In this study, for the noisy gluonic matrix elements, $\omega$ is the only input feature in the parametrization of the ML. This paper aims to lay the groundwork for ML applications on future precise lattice data. Here, we examine how the LQCD data can be extrapolated to an $\omega$-region beyond the range of the currently available data, enabling more accurate determination of PDFs while minimizing unwanted contamination using the generative models of ML.

### v2

@@ I Introduction
To achieve the goal of calculating momentum fraction, $x$-dependent nonperturbative structure functions and PDFs from the first-principles lattice QCD (LQCD) calculations, there have been several proposals, such as the path-integral formulation of the deep-inelastic scattering hadronic tensor Liu and Dong (1994); Liu (2000), the operator product expansion Detmold and Lin (2006), current-current correlations Braun and Müller (2008), the Compton amplitude approach Horsley et al. (2012); Chambers et al. (2017), quasi-PDFs and large momentum effective theory (LaMET) Ji (2013, 2014), lattice cross-sections Ma and Qiu (2018a, b); Sufian et al. (2019), and pseudo-PDFs Radyushkin (2017). While LQCD calculations have made significant computational achievements in calculating PDFs (see the recent reviews Ji et al. (2021a); Constantinou et al. (2022) and the references therein), one defining challenge is that the bilocal light-cone correlators that are necessary to determine the PDFs cannot be evaluated directly on the Euclidean lattice. Quasi-PDFs framework Ji (2013) circumvents this drawback by calculating equal-time Euclidean nonlocal matrix elements with hadron states at non-zero momentum, $p_{z}$. The corresponding quasi-PDFs can be matched to the light-cone PDFs when the hadron momentum is large, by applying the LaMET formalism Ji (2014). On the other hand, the pseudo-PDFs framework uses short-distance factorization and can be perturbatively matched to the light-cone PDF. To implement the QCD short-distance factorization with controllable power corrections and because the renormalons in the renormalization of the LQCD matrix elements at large distances become significantly important Braun et al. (2019), it is desirable to use only the short-distance matrix elements. However, one needs LQCD matrix elements at large Ioffe-time $\omega\equiv-z\cdot p$ Gribov et al. (1965); Ioffe (1969); Braun et al. (1995) while keeping $z$ small and making the hadron boost $p_{z}$ along the $z$-direction as large as possible to facilitate the inverse problem in determining PDFs from a limited $\omega$-range data. The problem of using large-$z$ matrix elements has been found in a recent calculation Bhat et al. (2022) with the implementation of the 2-loop matching Li et al. (2021) as the higher-twist contribution can become significant for $z\gtrsim 0.35$ fm (see for other findings in Karthik and Sufian (2021); Ji (2022); Su et al. (2022)). An alternative proposal is to use the pseudo-PDFs ratio scheme Radyushkin (2017) only at small $z$ in hybrid- and self-renormalization schemes Ji et al. (2021b); Huo et al. (2021), treating short and long-distance scales separately as well as using model-dependent extrapolations of lattice correlation functions at large $\omega$ to facilitate the determination of PDFs within the LaMET framework.
@@
In this work, we train the generative machine learning (ML) algorithms using short-distance LQCD correlation functions and focus on the generation of the LQCD matrix elements at large $\omega$ which can be used to determine the polarized gluon $x\Delta g(x)$ and the unpolarized gluon $xg(x)$ PDFs. The principle behind training ML algorithms with short-distance correlation functions is based on the ability of the ML to learn effectively from the data points where higher-twist contamination is minimal. Therefore, it is expected that the generation of the renormalized matrix elements at large $\omega$ will also involve small contamination from higher-twist effects. Moreover, the ML algorithms predicting the LQCD matrix element at large $\omega$ for the larger values of $p_{z}$ will not only be extremely useful for handling the inverse problem of reconstructing the PDFs but also be particularly helpful in future calculation of quark and gluon PDFs within the LaMET fromalism Ji (2014).
@@
We note that there have been a few calculations of the unpolarized gluon PDFs in the nucleon, pion, and kaon Fan et al. (2021); Fan and Lin (2021); Fan et al. (2022); Salas-Chavira et al. (2021); Khan et al. (2021); Delmar et al. (2023); Good et al. (2023); Good and Lin (2024); Good et al. (2024), as well as matrix element calculations toward the gluon helicity Ioffe-time distribution Egerer et al. (2022), and the first LQCD determination of the gluon helicity PDF Khan et al. (2023). In our work, we will show that determining the light-cone quark and gluon helicity PDFs require LQCD data in the limit of $p_{z}\to\infty$. We will demonstrate how physics-informed ML can help attain this limit and enable us to determine helicity PDFs, particularly the gluon helicity PDF. We will use published LQCD matrix elements of the unpolarized and polarized gluon correlation functions Balitsky et al. (2020, 2022) from Refs. Khan et al. (2021); Egerer et al. (2022); Khan et al. (2023). The essential goals of this work are as follows:
@@
The paper is organized as follows. In Sec. II, we first discuss the necessity of extrapolating LQCD data outside the available range of $\omega$ and discuss how this extrapolation has been performed previously. This section reveals the risk of using model-dependent and overly constrained PDF ansatz and motivates the need for ML to generate data at large $\omega$ where LQCD data are not available. In Sec. III, we discuss the theoretical essence of LQCD data at large momentum and the prospect of generative ML models to overcome the above-mentioned problems. We discuss a prescription to eliminate the $p_{z}$-dependence in the lattice data of quark and gluon helicity correlation functions. In Secs. IV and V, we explore three different generative ML algorithms to determine the polarized and unpolarized gluon Ioffe-time distributions Braun et al. (1995) beyond the reach of present-day LQCD calculations. In Secs. VI and VII, we determine $x\Delta g(x)$ and $xg(x)$ distributions using neural network on the ML-generated Ioffe time distributions and highlight some of the important implications of this work. Sec. VIII contains our concluding remarks and outlook.
@@
In general, LQCD calculation of matrix elements at hadron momentum $p_{z}>2~{}{\rm to}~{}3$ GeV becomes exponentially challenging due to poor signal-to-noise ratio. Specifically, achieving precise results for the gluonic matrix elements at a lighter pion mass and $p_{z}>2$ GeV will be very challenging in the near future LQCD calculations as it involves Wick contractions associated with the lattice calculations of disconnected insertions. The state-of-art LQCD calculations of the unpolarized gluonic matrix elements or the Ioffe-time distribution (ITD) have provided data up to $\omega_{\rm max}\approx 7$ Khan et al. (2021); Fan et al. (2022) or even smaller Delmar et al. (2023), beyond which the data is consistent with zero due to large noise. One way to facilitate the extraction of PDFs from the lattice data in a wide range of $\omega$, is to perform LQCD calculation at large spacelike separations at a fixed $p_{z}$. However, this poses the risk of uncontrollable power corrections, and the factorization of the LQCD matrix elements into nonpertubative PDF and perturbative coefficients might not be sensible. Additionally, at a fixed $p_{z}$, lattice data becomes much noisier as $z$ increases. Therefore, when one parameterizes the PDFs using some model-constrained functional forms and includes possible higher-twist contributions such as $\mathcal{O}(\Lambda_{\rm QCD}^{2}z^{2})$ or some higher order $\mathcal{O}(z^{n})$ terms in parameterizing the data, the noisy data may not allow for proper isolation of the higher-twist contributions from the leading twist-2 contribution (for example, see Fig. 5 in Edwards et al. (2023) at $z=0.75$ fm and $p_{z}\geq 1.67$ GeV).
@@
In addition, constrained by a limited range of $\omega$, the available lattice ITDs are inherently sensitive to only the first few moments, as shown in Mankiewicz and Weigl (1996) and reiterated in Gao et al. (2020); Sufian et al. (2021); Su et al. (2022). Lattice calculations using model-dependent parametrization of PDFs utilize the correlation between the first two or three accessible moments to model the missing information beyond the available $\omega_{\rm max}$ and attempt to extract PDF in the full $x$ range. In addition, for currently available LQCD calculations in a limited $\omega$ range, these functional forms can be biased, leading to an unreliable $\chi^{2}/{\rm d.o.f.}$, and underestimation of uncertainties. For example, the parametrization $x^{\alpha}(1-x)^{\beta}$ for the LQCD determination of $xg(x)$ distribution led to a diverging PDF in Fan et al. (2021) and a converging PDF in Khan et al. (2021) at small $x$. However, none of these lattice ITDs reach the Regge region or have much sensitivity to the small-$x$ physics Sufian et al. (2021). In Khan et al. (2021), $\alpha\geq 0$ constraint was imposed in a Bayesian fit, motivated by a phenomenological analysis in Sufian et al. (2021). Otherwise, it would have resulted in a diverging PDF as in Fan et al. (2021). A similar observation can be made for $xg(x)$ extractions in recent lattice calculations Fan et al. (2018); Fan and Lin (2021); Fan et al. (2022); Salas-Chavira et al. (2021); Delmar et al. (2023); Good et al. (2023) where all these lattice ITDs are limited in ITD the range of $\omega\lesssim 6$. It is therefore desirable to determine LQCD matrix elements at large $\omega$, while maintaining controllable power corrections. Additionally, it is important not to rely on simple ansatz fits of PDFs using data in a limited $\omega$ range as the resulting PDFs can depend heavily on the chosen functional forms, priors, and other strong constraints used in the fits. None of the existing LQCD data is sensitive to either very large $x$ or very small $x$ PDFs. It is not evident why a product of $x^{\alpha}$ (governed by the Regge behavior Regge (1959)) and $(1-x)^{\beta}$ (governed by the perturbative QCD counting rule Brodsky and Farrar (1973); Lepage and Brodsky (1980); Brodsky et al. (1995)) should describe PDFs in the mid-$x$ region, where the lattice data is more sensitive to.
@@
As an illustration of the problem, by reconstructing the unpolarized gluon ITD from NNPDF distribution Ball et al. (2017), it can be shown that the NNPDF ITD has significantly smaller uncertainty across the entire $\omega$ range compared to the lattice ITD presented in Khan et al. (2021). However, the large-$x$ PDF determined from the LQCD data has much smaller uncertainty than the NNPDF distribution. Although the Fourier transform of $\tilde{\omega}=\frac{1}{x}$ Ma and Qiu (2018b); Miller and Brodsky (2020) and precise lattice data at small $\omega$ can be helpful to constrain PDF in the large-$x$ region, such a dramatic reduction of the uncertainty in the large-$x$ PDFs is not expected. This possibility of underestimating the uncertainty in the lattice calculation of PDFs has been elaborated in Fig. 1, where the lattice and the phenomenological ITDs $\mathcal{I}_{g}(\omega,\mu)$ are chosen to be at $\mu=2$ GeV matching scale in the $\overline{\rm{MS}}$ scheme. Keeping in mind that this is not a one-to-one comparison because of the different types of observables used in the determination of PDFs, the reconstructed PDF from the lattice data using a Bayesian fit with the functional form of PDF $x^{\alpha}(1-x)^{\beta}$ (with imposed constraint $\alpha\geq 0$) in Fig. 1 has uncertainty smaller than the NNPDF determination in the whole $x\geq 0.7$ region. Therefore, it can be misleading to obtain a much more precise PDF from noisier lattice data in a limited $\omega$ range. The reason for this unrealistic precision in Khan et al. (2021) stems from two constraints:
@@
  • Bayesian fits using Jacobi polynomial basis Atashbar Tehrani and Khorramian (2007); Khorramian et al. (2009); Taghavi-Shahri et al. (2016) with a constraint $\alpha\geq 0$.
A more direct comparison using the same lattice data with or without using constrained functional forms in the determination of gluon helicity PDFs has been illustrated in the lower panel in Fig. 1 (from Ref. Khan et al. (2023)) which shows the severity of underestimating uncertainties when constrained PDF ansatz is used.
@@
Figure 1: Upper left panel: Ioffe-time distribution after the implementation of the perturbative matching kernel on the lattice reduced pseudo-ITD calculated for a $2$-parameter model fit, in the $\overline{\rm{MS}}$ renormalization scheme at 2 GeV in Ref. Khan et al. (2021). The $z=a-6a$ denotes the length of the field separation as a function of the lattice spacing $a$. The Ioffe-time distribution of the NNPDF distribution is normalized to $1$ at $\omega=0$ for a comparison with the LQCD matrix elements. Upper right panel: unpolarized gluon PDF (cyan band) extracted from the lattice data in Ref. Khan et al. (2021) using the $2$-parameter model fit $x^{\alpha}(1-x)^{\beta}$. We compare our results to the gluon PDF extracted from the NNPDF3.1 global fit Ball, Richard D. et al. (2017). Lower panel: a similar example is given for the gluon helicity PDF from Khan et al. (2023). The cyan band is determined directly from the lattice QCD data without assuming any functional form for the PDF ansatz, while the other bands with smaller uncertainties result from different model fits used to extract the $x\Delta g(x)$ distribution. The 2-parameter fit is done using the fit-expression $x^{\alpha}(1-x)^{\beta}$, and subsequently by introducing additional parameters for 3 and 4 parameter fits, namely $\rho$ and $\gamma$, resulting in the fit-expression $x^{\alpha}(1-x)^{\beta}(1+\rho\sqrt{x}+\gamma x)$.
@@
Building on the findings presented in Khan et al. (2023), this section briefly outlines the prescription for removing contamination terms in the Euclidean matrix elements that may obstruct the extraction of light-cone quark and gluon helicity PDFs.
To determine $\Delta g(x)$, we calculate matrix elements of the gluon field strength tensor $G_{\mu\nu}$ and its dual $\widetilde{G}_{\lambda\beta}=(1/2)\epsilon_{\lambda\beta\rho\gamma}G^{\rho\gamma}$ separated by a spatial Wilson line $W[z,0]$ Ji (2013); Balitsky et al. (2022),
@@
where $z$ is the separation between the gluon fields, $p$ is the nucleon four-momentum, and $s$ is the nucleon polarization. The combination that can be used to determine the gluon helicity correlation with the least number of contamination terms is Balitsky et al. (2022)
@@
$i,j=x,y$ being perpendicular to the nucleon boost in the $z$-direction, $p=\{p_{0},0_{\perp},p_{z}\}$. Utilizing the multiplicative renormalizability of the quasi-PDF matrix elements Izubuchi et al. (2018); Ji et al. (2018); Zhang et al. (2019); Li et al. (2019), one can use the pseudo-PDFs ratio method Radyushkin (2017) and obtain the renormalization group invariant reduced pseudo-ITD, $\Delta\mathfrak{{M}}$ Balitsky et al. (2022):
@@
where, $m_{p}$ is the nucleon mass, ${\cal M}_{00}(z,p_{z})\equiv[M_{0i;i0}(z,p_{z})+M_{ji;ij}(z,p_{z})]$ is the spin-averaged matrix element related to the unpolarized gluon correlation Balitsky et al. (2020); Khan et al. (2021) and the factor $1/Z_{\rm L}(z/a_{L})$ is determined in Balitsky et al. (2022) to cancel the UV logarithmic vertex anomalous dimension of $\Delta{\cal M}_{00}$.
$\Delta\mathfrak{{M}}$ can be expressed in terms of invariant amplitudes, $\Delta\mathcal{M}_{sp}^{(+)}$ and $\Delta{\mathcal{M}}_{pp}$ Balitsky et al. (2022),
@@
where the subscript $lc$ denotes taking the light-cone limit $z=z_{-}$ and applying the UV renormalization resulting in the dependence on the factorization scale $\mu$. Note that light-cone ITD in Eq. (5) does not contain the $m_{p}^{2}/p_{z}^{2}$ suppressed term as appeared in Eq. (4). An alternative expression of $\Delta\mathcal{M}_{00}(z,p_{z})$ shows that this matrix element is nonvanishing at $p_{z}=0$ and one can remove the $\mathcal{O}(\omega)$ contamination following the method describing in Egerer et al. (2022). However, the residual higher-order contamination can become significant at large $\omega$. In order to eliminate the $(m_{p}^{2}/p_{z}^{2})\omega\Delta{\mathcal{M}}_{pp}$ contribution, we multiply Eq. (4) by the corresponding lattice squared-momentum $p_{k}^{2}$ and obtain Khan et al. (2023)
@@
Following the derivation in Saalfeld et al. (1998), it can be shown that in a fixed gauge the nonlocal spatial matrix element in Eq. (2) reduces to $\Delta G\rightarrow\left(\vec{E}\times\vec{A_{\perp}}\right)^{3}$ in the infinite momentum frame, consistent with the findings in Ji et al. (2013); Hatta et al. (2014), where $\vec{A}_{\perp}$ is the transverse part of the gauge potential.
As another example, a contamination term is also present in the LQCD matrix elements associated with the quark helicity PDF Edwards et al. (2023). Following the same notations as in Edwards et al. (2023), the Euclidean matrix elements needed for the calculation of quark helicity PDF can be written as
@@
Now, the essence of the ML algorithms to utilize the relation in Eq. (7) can be understood as the following. The immediate challenge of implementing Eq. (7) is that the amplitudes $\Delta{\mathcal{M}}_{sp}^{(+)}$ and $\Delta{\mathcal{M}}_{pp}$ are nonperturbative functions and their functional forms are unknown. It has been shown in Khan et al. (2023) that in the absence of a theoretically known functional form of these nonperturbative amplitudes, the moments expansion fails to describe the data, and one needs to resort to ML to circumvent the ignorance of the functional forms of these amplitudes. Therefore, to parameterize these unknown nonperturbative amplitudes, one can use some ML approaches constrained by Eq. (7) which we refer to as the “physics-informed” ML. Note that, as long as the equality in Eq. (7) holds to a good approximation, the lattice correlation functions will be dominated by the leading-twist contribution and can be used to extract the light-cone $x\Delta g(x)$. As a remarkable outcome, we will see in Sec. IV.1 that this relation is beneficial in exposing possible higher-twist contaminations in the gluon correlation functions at $z\gtrsim 0.36$ fm. Therefore, if ML approaches can be developed and trained on the lattice matrix elements at small $z$, where higher-twist contamination is minimal, and then used to generate the ITD at large $\omega$, the resulting extrapolated data may exhibit reduced higher-twist contamination even at larger values of $\omega$ for a fixed $p_{z}$. The ideal scenario is when ample and precise lattice data points are available, allowing ML to determine the correlations between different $z$ and $p_{z}$ datasets and estimate higher-twist contaminations. However, the gluon matrix elements do not allow us to isolate the $z$-dependence across different $p_{z}$ datasets. For example, due to the large uncertainties in the data points, we were unable to isolate the $p_{z}$-dependence in the unpolarized gluon matrix elements.This will be a subject of future study with more precise LQCD matrix elements. In this study, for the noisy gluonic matrix elements, $\omega$ is the only input feature in the parametrization of the ML. This does not imply that different $p_{z}$ datasets are independent; rather, it indicates that we are unable to separately disentangle the $z$- and $p_{z}$-dependence from the existing datasets. This paper aims to lay the groundwork for ML applications on future precise lattice data. Here, we examine how the LQCD data can be extrapolated to an $\omega$-region beyond the range of the currently available data, enabling more accurate determination of PDFs while minimizing unwanted contamination using the generative models of ML.